# and Qian Liu

### v1

@@ III-B.1 Long-timescale EM domain Channel Estimation
Given the specific scattering cluster characteristics described in (4), the angles of the multi-path rays within each scattering cluster exhibit only limited fluctuations around the cluster core. This observation implies that real-time radiation patterns adaptation based on instantaneous CSI is unnecessary. Instead, aligning radiation patterns with the nominal angles of scattering clusters is sufficient to maintain robust communication performance, while significantly reducing channel estimation overhead and beamforming design complexity. Therefore, over the long-timescale, our objective is to estimate the nominal angles and average channel gains of the scattering clusters, which serve as the foundation for radiation beamforming optimization.
@@
Specifically, we employ the Capon algorithm to construct the angular power spectrum (APS) based on the channel covariance matrix, enabling precise spatial domain analysis [29]. To illustrate the channel estimation process, we consider the $k$-th UE as an example. By accumulating spatial-channel samples over the medium-timescale, the average channel covariance matrix at the $t_{\mathrm{L}}$-th super-frame is given by
$\mathbf{R}_{k}^{(t_{\mathrm{L}})}=\frac{1}{t_{\mathrm{L}}T_{\mathrm{M}}}\sum_{t_{\mathrm{l}}=1}^{t_{\mathrm{L}}}\sum_{t_{\mathrm{M}}=1}^{T_{\mathrm{M}}}\widetilde{\mathbf{h}}^{(t)}_{\mathrm{S},k}(\widetilde{\mathbf{h}}_{\mathrm{S},k}^{(t)})^{H},~{}\forall k,t_{\mathrm{L}},$ (17)
@@
where $t=(t_{\mathrm{l}}-1)T_{\mathrm{M}}+t_{\mathrm{M}}$ is defined for brevity and $\widetilde{\mathbf{h}}_{\mathrm{S},k}$ represents the estimated channel excluding the influence of the radiation patterns, given by
$\widetilde{\mathbf{h}}_{\mathrm{S},k}=\sum_{l=1}^{L}\alpha_{k,l}\mathbf{a}_{k,l},~{}\forall k.$ (18)
Then, the APS can be derived as
$\rho^{(t_{\mathrm{L}})}_{k}(\theta)=\frac{1}{\mathbf{e}^{H}(\theta)(\mathbf{R}_{k}^{(t_{\mathrm{L}})})^{-1}\mathbf{e}(\theta)},~{}\forall k,t_{\mathrm{L}},$ (19)
with
$\mathbf{e}(\theta)\triangleq[1,e^{-\jmath\frac{2\pi}{\lambda}d\sin\theta},\ldots,e^{-\jmath\frac{2\pi}{\lambda}(N_{\mathrm{t}}-1)d\sin\theta}]^{T}\in\mathbb{C}^{N_{\mathrm{t}}}.$ (20)
@@
Since the APS reflects the power distribution across the angular domain, each peak in the APS physically represents the presence of a distinct scattering cluster in the propagation environment. Thus, the number of the peaks in $\rho^{(t_{\mathrm{L}})}_{k}(\theta)$ that exceed a specified threshold directly provides the estimated number of scattering clusters $C$. For each of these peaks, the corresponding angle value and power value indicate the estimated nominal angle $\overline{\theta}_{k,c}$ and the average channel gain $\alpha_{k,c}$ of the corresponding scattering cluster, respectively. In conclusion, referring to (12), we can construct the EM domain channel, which serves as the foundation for the long-timescale EM domain radiation beamformer design, which yields
$\mathbf{h}_{\mathrm{C},k}=\overline{\mathbf{H}}_{\mathrm{EM},k}\overline{\mathbf{h}}_{\mathrm{S},k}\in\mathbb{C}^{MN_{\mathrm{t}}},$ (21)
@@
where the term $\overline{\mathbf{h}}_{\mathrm{S},k}\in\mathbb{C}^{CN_{\mathrm{t}}}$ encapsulates the spatial domain channel characteristics of $C$ scattering cluster cores and $\mathbf{\overline{H}}_{\mathrm{EM},k}\in\{0,1\}^{MN_{\mathrm{t}}\times CN_{\mathrm{t}}}$ indicates the angle information of the scattering clusters.
@@ III-B.2 Long-timescale Radiation Beamformer Design
Notably, over the long-timescale, the EM domain radiation beamformer $\mathbf{F}_{\mathrm{EM}}$ mainly aims to establish a basic pattern alignment with the scattering cluster core. This optimization requires only the channel information of the scattering cluster core, i.e., the estimated nominal angles and average channel gains of the scattering clusters. Consequently, we substitute the instantaneous multi-path channel $\mathbf{h}_{k}$ in the SINR expression (14) and the sum-rate formulation (15) with the scattering cluster core channel $\mathbf{h}_{\mathrm{C},k}$, defining the long-timescale sum-rate $\overline{R}_{k}$ as our optimization objective. Given a fixed analog beamformer $\mathbf{F}_{\mathrm{RF}}$ and digital beamformer $\mathbf{F}_{\mathrm{BB}}$, the long-timescale EM domain radiation beamforming design problem is then formulated as
$\max_{\mathbf{F}_{\mathrm{EM}}}$ $\ \ \sum^{K}_{k=1}\overline{R}_{k}$ (22a)
@@
where $\overline{\mathbf{F}}_{\mathrm{pat}}\triangleq\mathbf{I}_{N_{\mathrm{t}}}\otimes\mathbf{F}_{\mathrm{pat}}\in\mathbb{R}^{MN_{\mathrm{t}}\times PN_{\mathrm{t}}}$ with $\mathbf{F}_{\mathrm{pat}}\in\mathbb{R}^{M\times P}$ representing a predefined radiation pattern dictionary that contains the radiation gains of $P$ candidate patterns over $M$ spatial sampling angles. The matrix $\mathbf{S}_{\mathrm{EM}}\triangleq\mathrm{blkdiag}\{\mathbf{s}_{\mathrm{EM},1},\mathbf{s}_{\mathrm{EM},2},\ldots,\mathbf{s}_{\mathrm{EM},N_{\mathrm{t}}}\}\in\{0,1\}^{PN_{\mathrm{t}}\times N_{\mathrm{t}}}$ denotes the EM domain pattern selection matrix, where each $\mathbf{s}_{\mathrm{EM},n}\in\{0,1\}^{P}$ selects a single optimal pattern from the dictionary for the $n$-th antenna, i.e., $\|\mathbf{s}_{\mathrm{EM},n}\|_{1}=1$. Based on this formulation, the original problem in (22) is equivalently transformed into a more practical optimization problem focused on optimizing the EM-domain pattern selection matrix $\mathbf{S}_{\mathrm{EM}}$, given by
$\max_{\mathbf{S}_{\mathrm{EM}}}$ $\ \ \sum^{K}_{k=1}\overline{R}_{k}$ (24a)
@@ III-C.1 Medium-timescale RF Domain Channel Estimation
Initially, the RF domain channel to be estimated is expressed as
$\mathbf{h}_{\mathrm{RF},k}\triangleq\mathbf{F}_{\mathrm{EM}}^{H}\mathbf{\overline{H}}_{\mathrm{EM},k}\overline{\mathbf{h}}_{\mathrm{S},k},~{}\forall k,$ (37)
@@
where $\mathbf{F}_{\mathrm{EM}}^{H}\mathbf{\overline{H}}_{\mathrm{EM},k}$ reflects the radiation gains of the reconfigurable antennas based on the fixed radiation beamformer determined in the long-timescale, and it is apparent that this term is coupled with the spatial channel $\overline{\mathbf{h}}_{\mathrm{S},k}$. During the training phase, all UEs simultaneously transmit pilot signals on distinct frequency tones to the BS in the uplink. Taking the $k$-th UE as an example, the received signal at the BS is given by
$\mathbf{y}_{k}$ $=\sqrt{P_{\mathrm{t}}}\mathbf{W}_{k}^{H}\mathbf{h}_{\mathrm{RF},k}+\mathbf{W}_{k}^{H}\mathbf{n}_{k}$ (38)
$=\sqrt{P_{\mathrm{t}}}\mathbf{W}_{k}^{H}\mathbf{A}_{\mathrm{D}}\mathbf{z}_{k}+\mathbf{W}_{k}^{H}\mathbf{n}_{k}$
$=\mathbf{Q}_{k}\mathbf{z}_{k}+\mathbf{W}_{k}^{H}\mathbf{n}_{k},~{}\forall k,$
where $\mathbf{W}_{k}\in\mathbb{C}^{N_{\mathrm{t}}\times N_{\mathrm{Beam}}}$ is the training beamformer with $N_{\mathrm{Beam}}$ denoting the number of training beams, $\mathbf{A}_{\mathrm{D}}\in\mathbb{C}^{N_{\mathrm{t}}\times M}$ is the refined dictionary matrix incorporating the effects of pattern-reconfigurable antennas, $\mathbf{z}_{k}$ is a sparse vector with non-zero elements corresponding to the spatial angles, and $\mathbf{Q}_{k}\triangleq\sqrt{P_{\mathrm{t}}}\mathbf{W}_{k}^{H}\mathbf{A}_{\mathrm{D}}$ is the sensing matrix.
@@
The conventional dictionary matrix consists of a set of steering vectors corresponding to uniformly distributed candidate angles, given by
$\mathbf{E}=[\mathbf{e}(\theta_{1}),\mathbf{e}(\theta_{2}),\ldots,\mathbf{e}(\theta_{M})]\in\mathbb{C}^{N_{\mathrm{t}}\times M},$ (39)
@@
where each column $\mathbf{e}(\theta_{m})$ represents a steering vector pointing toward the angle $\theta_{m}=\pi m/M,~{}m=1,2,\ldots,M$. To explicitly account for the influence of the reconfigurable antenna pattern on the channel representation, we refine the dictionary matrix as
$\mathbf{A}_{\mathrm{D}}(:,m)=\mathbf{F}_{\mathrm{EM}}(:,m)\odot\mathbf{e}(\theta_{m}),~{}\forall m,$ (40)
@@
where $\mathbf{F}_{\mathrm{EM}}(:,m)$ characterizes the radiation gains of the $N_{\mathrm{t}}$ antennas in the direction of $\theta_{m}$. This formulation effectively integrates both radiation pattern effects and angular information, enabling a more accurate estimation of spatial channel characteristics.
@@
$\min_{\mathbf{z}_{k}}$ $\ \ \ \|\mathbf{y}_{k}-\mathbf{Q}_{k}\mathbf{z}_{k}\|_{2}$ (41a)
@@
This problem is typically solved using the orthogonal matching pursuit (OMP) algorithm, which iteratively selects the most relevant dictionary elements to approximate the signal. The iterations continue until the residual error falls below a predefined threshold or the sparsity constraint is met. The implementation details of the OMP algorithm and the design of the training hybrid beamformer follow the principles outlined in [32], [33]. For brevity, these details are omitted here.

### v2

@@ III-B.1 Long-timescale EM domain Channel Estimation
Given the specific scattering cluster characteristics described in (7), the angles of the multi-path rays within each scattering cluster exhibit only limited fluctuations around the cluster core. This observation implies that real-time radiation patterns adaptation based on instantaneous CSI is unnecessary. Instead, aligning radiation patterns with the nominal angles of scattering clusters is sufficient to maintain robust communication performance, while significantly reducing channel estimation overhead and beamforming design complexity. Therefore, over the long-timescale, our objective is to estimate the nominal angles $\overline{\theta}_{k,c}$ and the mean large-scale channel gains $\alpha_{k,c}$ of the scattering clusters, which serve as the foundation for radiation beamforming optimization.
@@
Specifically, we employ the Capon algorithm to construct the angular power spectrum (APS) based on the channel covariance matrix, enabling the extraction of scattering cluster channel parameters [32], [33]. To illustrate the channel estimation process, we consider the $k$-th UE as an example. In the first super-frame, all the pattern-reconfigurable antennas are typically configured to radiate toward the boresight direction. Starting from the second super-frame, the scattering-cluster channel information is estimated by accumulating spatial channel samples collected over the medium-timescale of the preceding super-frames. Accordingly, the channel covariance matrix at the $t_{\mathrm{L}}$-th super-frame is given by
$\mathbf{R}_{k}^{(t_{\mathrm{L}})}=\frac{1}{t_{\mathrm{L}}T_{\mathrm{M}}}\sum_{t_{\mathrm{l}}=1}^{t_{\mathrm{L}}-1}\sum_{t_{\mathrm{M}}=1}^{T_{\mathrm{M}}}\widetilde{\mathbf{h}}^{(t)}_{\mathrm{S},k}(\widetilde{\mathbf{h}}_{\mathrm{S},k}^{(t)})^{H},~\forall k,t_{\mathrm{L}},$ (19)
@@
where $t=(t_{\mathrm{l}}-1)T_{\mathrm{M}}+t_{\mathrm{M}}$ is defined for brevity and $\widetilde{\mathbf{h}}_{\mathrm{S},k}\triangleq\sum_{l}^{L}\alpha_{k,l}e^{j\phi_{k,l}}\mathbf{a}_{k,l}$ represents the estimated channel excluding the influence of the radiation patterns. Then, the APS can be derived as
$\rho^{(t_{\mathrm{L}})}_{k}(\vartheta_{m})=\frac{1}{\mathbf{e}^{H}(\vartheta_{m})(\mathbf{R}_{k}^{(t_{\mathrm{L}})})^{-1}\mathbf{e}(\vartheta_{m})},~\forall k,t_{\mathrm{L}},$ (20)
@@
with $\mathbf{e}(\vartheta_{m})\!\triangleq\![1,e^{-\jmath\frac{2\pi}{\lambda}d\sin\vartheta_{m}},\ldots,e^{-\jmath\frac{2\pi}{\lambda}(N_{\mathrm{t}}-1)d\sin\vartheta_{m}}]^{T}\in\mathbb{C}^{N_{\mathrm{t}}}.$ Since the APS reflects the power distribution across the angular domain, each peak in the APS physically represents the presence of a distinct scattering cluster in the propagation environment. Thus, the number of the peaks in $\rho^{(t_{\mathrm{L}})}_{k}(\vartheta_{m})$ that exceed a specified threshold directly provides the estimated number of scattering clusters $C$. For each of these peaks, the corresponding angle value and power value indicate the estimated nominal angle $\overline{\theta}_{k,c}$ and the mean large-scale channel gain $\alpha_{k,c}$ of the corresponding scattering cluster, respectively. In conclusion, referring to (14), we can construct the channels of scattering clusters as
$\mathbf{h}_{\mathrm{C},k}=\mathbf{H}_{\mathrm{EM},\mathrm{C},k}\mathbf{h}_{\mathrm{S},\mathrm{C},k}\in\mathbb{C}^{MN_{\mathrm{t}}},~\forall k,$ (21)
@@
where $\mathbf{H}_{\mathrm{EM},\mathrm{C},k}\triangleq[\mathbf{H}_{\mathrm{EM},\mathrm{C},k,1},\mathbf{H}_{\mathrm{EM},\mathrm{C},k,2},\ldots,\mathbf{H}_{\mathrm{EM},\mathrm{C},k,C}]\in\{0,1\}^{MN_{\mathrm{t}}\times CN_{\mathrm{t}}}$ indicates the EM domain angle information and $\mathbf{h}_{\mathrm{S},\mathrm{C},k}\triangleq[\mathbf{h}_{\mathrm{S},\mathrm{C},k,1}^{T},\mathbf{h}_{\mathrm{S},\mathrm{C},k,2}^{T},\ldots,\mathbf{h}_{\mathrm{S},\mathrm{C},k,C}^{T}]^{T}\in\mathbb{C}^{CN_{\mathrm{t}}}$ encapsulates the spatial domain channel characteristics of $C$ scattering cluster cores. It serve as the foundation for the long-timescale EM domain radiation beamformer design.
@@ III-B.2 Long-timescale Radiation Beamformer Design
Notably, over the long-timescale, the EM domain radiation beamformer $\mathbf{F}_{\mathrm{EM}}$ mainly aims to establish a basic pattern alignment with the scattering cluster core. This optimization requires only the channel information of the scattering cluster core, i.e., the estimated nominal angles and average channel gains of the scattering clusters. Consequently, we substitute the instantaneous multi-path channel $\mathbf{h}_{k}$ in the SINR expression (16) and the sum-rate formulation (17) with the scattering cluster core channel $\mathbf{h}_{\mathrm{C},k}$, and define the long-timescale sum-rate $R_{\mathrm{C},k}$ as our optimization objective. Given a fixed analog beamformer $\mathbf{F}_{\mathrm{RF}}$ and digital beamformer $\mathbf{F}_{\mathrm{BB}}$, the long-timescale EM domain radiation beamforming design problem is then formulated as
$\max_{\mathbf{F}_{\mathrm{EM}}}$ $\ \ \sum^{K}_{k=1}R_{\mathrm{C},k}$ (22a)
@@
where $\overline{\mathbf{F}}_{\mathrm{pat}}\triangleq\mathbf{I}_{N_{\mathrm{t}}}\otimes\mathbf{F}_{\mathrm{pat}}\in\mathbb{R}^{MN_{\mathrm{t}}\times PN_{\mathrm{t}}}$ with $\mathbf{F}_{\mathrm{pat}}\in\mathbb{R}^{M\times P}$ representing a predefined radiation pattern dictionary that contains the radiation gains of $P$ candidate patterns over $M$ spatial sampling angles. The matrix $\mathbf{S}_{\mathrm{EM}}\triangleq\mathrm{blkdiag}\{\mathbf{s}_{\mathrm{EM},1},\mathbf{s}_{\mathrm{EM},2},\ldots,\mathbf{s}_{\mathrm{EM},N_{\mathrm{t}}}\}\in\{0,1\}^{PN_{\mathrm{t}}\times N_{\mathrm{t}}}$ denotes the EM domain pattern selection matrix, where each sub-vector $\mathbf{s}_{\mathrm{EM},n}\in\{0,1\}^{P}$ selects a single optimal pattern from the dictionary for the $n$-th antenna, i.e., $\|\mathbf{s}_{\mathrm{EM},n}\|_{1}=1$. Based on this formulation, the original problem in (22) is equivalently transformed into a more practical optimization problem focused on optimizing the EM domain pattern selection matrix $\mathbf{S}_{\mathrm{EM}}$, given by
$\max_{\mathbf{S}_{\mathrm{EM}}}$ $\ \ \sum^{K}_{k=1}R_{\mathrm{C},k}$ (24a)
@@ III-C.1 Medium-timescale RF Domain Channel Estimation
Initially, the RF domain channel at each frame to be estimated is expressed as
$\mathbf{h}_{\mathrm{RF},k}\triangleq\mathbf{F}_{\mathrm{EM}}^{H}\mathbf{H}_{\mathrm{EM},k}\mathbf{h}_{\mathrm{S},k},~\forall k,$ (37)
@@
where $\mathbf{F}_{\mathrm{EM}}^{H}\mathbf{H}_{\mathrm{EM},k}$ reflects the radiation gains of the reconfigurable antennas based on the fixed radiation beamformer determined in the long-timescale, and it is apparent that this term is coupled with the spatial channel $\mathbf{h}_{\mathrm{S},k}$. During the uplink training phase, we assume the $K$ UEs transmit pilot sequences over orthogonal time resources. Therefore, the BS can isolate each UE’s signal, making channel estimation for different UEs completely independent. Suppose that $I$ pilot symbols are allocated to each UE. Taking the $k$-th UE as an example, the corresponding received signal during the $i$-th pilot at the BS $\mathbf{y}_{k}[i]\in\mathbb{C}^{N_{\mathrm{RF}}}$ can be expressed as
$\mathbf{y}_{k}[i]$ $=\mathbf{W}[i]^{H}\mathbf{h}_{\mathrm{RF},k}x[i]+\mathbf{W}[i]^{H}\mathbf{n}[i],~\forall k,i,$ (38)
where $\mathbf{W}[i]\triangleq\mathbf{F}_{\mathrm{RF}}[i]\mathbf{F}_{\mathrm{BB}}\in\mathbb{C}^{N_{\mathrm{t}}\times N_{\mathrm{RF}}}$ is the training combining matrix, $x[i]=1$ represents the transmit pilot, $\mathbf{n}[i]\sim\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I}_{N_{\mathrm{t}}})$ denotes the AWGN. The analog combining matrix $\mathbf{F}_{\mathrm{RF}}$ consists of constant-modulus entries randomly drawn from the quantized phase set $\mathcal{F}\triangleq\{\frac{1}{\sqrt{N_{\mathrm{t}}}}e^{\jmath\frac{2\pi b}{2^{B}}}|b=0,1,\ldots,2^{B}-1\}$ and the digital combining matrix is chosen as $\mathbf{F}_{\mathrm{BB}}\triangleq\mathbf{I}_{N_{\mathrm{RF}}}$ for simplicity. Collecting all $I$ pilot observations, the complete received pilot sequence at the BS for the $k$-th UE can be formulated as
$\mathbf{y}_{k}$ $=\mathbf{W}^{H}\mathbf{h}_{\mathrm{RF},k}+\mathbf{n},~\forall k,$ (39)
where
$\mathbf{y}_{k}$ $\triangleq[\mathbf{y}_{k}[1]^{T},\mathbf{y}_{k}[2]^{T},\ldots,\mathbf{y}_{k}[I]^{T}]^{T}\in\mathbb{C}^{N_{\mathrm{RF}}I},$ (40)
$\mathbf{W}$ $\triangleq[\mathbf{W}[1],\mathbf{W}[2],\ldots,\mathbf{W}[I]]\in\mathbb{C}^{N_{\mathrm{t}}\times N_{\mathrm{RF}}I},$
$\mathbf{n}$ $\triangleq[(\mathbf{W}[1]^{H}\mathbf{n}[1])^{T},(\mathbf{W}[2]^{H}\mathbf{n}[2])^{T},$
$\quad\quad\quad\quad\quad\quad\ldots,(\mathbf{W}[I]^{H}\mathbf{n}[I])^{T}]^{T}\in\mathbb{C}^{N_{\mathrm{RF}}I}.$
Considering that the number of RF chains is limited and excessive training overhead must be avoided, the dimension of the received pilot sequence is typically much smaller than the number of BS antennas, i.e., $N_{\mathrm{RF}}I\ll N_{\mathrm{t}}$. As a result, directly estimating the high-dimensional RF domain channel $\mathbf{h}_{\mathrm{RF},k}$ from the low dimensional observation $\mathbf{y}_{k}$ becomes highly challenging. Fortunately, the scattering cluster channel exhibits inherent sparsity in the angular domain. Leveraging this property, we adopt a CS–based channel estimation approach, under which the received signal in (39) can be equivalently represented as
$\mathbf{y}_{k}$ $=\mathbf{W}^{H}\mathbf{A}_{\mathrm{D}}\mathbf{z}_{k}+\mathbf{n}$ (41)
$=\mathbf{Q}\mathbf{z}_{k}+\mathbf{n},~\forall k,$
where $\mathbf{A}_{\mathrm{D}}\in\mathbb{C}^{N_{\mathrm{t}}\times M}$ is the refined dictionary matrix incorporating the effects of pattern-reconfigurable antennas, $\mathbf{z}_{k}\in\mathbb{C}^{M}$ is a sparse angular-domain coefficient vector whose nonzero entries correspond to spatial angles of the rays in the clusters, and $\mathbf{Q}\triangleq\mathbf{W}^{H}\mathbf{A}_{\mathrm{D}}$ is the sensing matrix, which is required to exhibit low mutual coherence and approximate Restricted Isometry Property (RIP) to enable accurate sparse recovery. Under such conditions, a CS-based method only requires the total number of effective measurements to exceed the number of propagation paths i.e., $N_{\mathrm{RF}}I\geq L$, to guarantee that $\mathbf{z}_{k}$ can be uniquely identified.
@@
We subsequently construct a refined dictionary matrix that explicitly embeds the radiation gain of pattern-reconfigurable antennas, yielding a more precise characterization of the spatial-domain channel. The conventional dictionary matrix consists of a set of steering vectors corresponding to uniformly distributed candidate angles, given by
$\mathbf{E}=[\mathbf{e}(\vartheta_{1}),\mathbf{e}(\vartheta_{2}),\ldots,\mathbf{e}(\vartheta_{M})]\in\mathbb{C}^{N_{\mathrm{t}}\times M},$ (42)
@@
where each column $\mathbf{e}(\vartheta_{m})\in\mathbb{C}^{N_{\mathrm{t}}}$ represents a steering vector pointing toward the angle $\vartheta_{m}=-\frac{\pi}{2}+\frac{\pi m}{M},~m=1,2,\ldots,M$. To incorporate the influence of the radiation pattern on the channel representation, we refine the dictionary matrix as
$\mathbf{A}_{\mathrm{D}}(:,m)=\overline{\mathbf{F}}_{\mathrm{EM}}(:,m)\odot\mathbf{e}(\vartheta_{m}),~\forall m,$ (43)
@@
where $\overline{\mathbf{F}}_{\mathrm{EM}}\triangleq[\mathbf{f}_{\mathrm{EM},1},\mathbf{f}_{\mathrm{EM},2},\ldots,\mathbf{f}_{\mathrm{EM},N_{\mathrm{t}}}]^{T}\in\mathbb{R}^{N_{\mathrm{t}}\times M}$ stores the radiation patterns of all antennas configured within the current super-frame, and $\overline{\mathbf{F}}_{\mathrm{EM}}(:,m)$ denotes the radiation gains of the $N_{\mathrm{t}}$ antennas in the direction of $\vartheta_{m}$. This formulation effectively integrates both radiation pattern effects and angular information, enabling a more accurate estimation of spatial channel characteristics. Moreover, to prevent directional bias and ensure uniform scaling across all angular atoms, each column of the refined dictionary is normalized as
$\mathbf{A}_{\mathrm{D}}(:,m)=\frac{\mathbf{A}_{\mathrm{D}}(:,m)}{\|\mathbf{A}_{\mathrm{D}}(:,m)\|},~\forall m.$ (44)
@@
$\min_{\mathbf{z}_{k}}$ $\ \ \ \|\mathbf{y}_{k}-\mathbf{Q}\mathbf{z}_{k}\|_{2}$ (45a)
@@
This problem is typically solved using the orthogonal matching pursuit (OMP) algorithm, which iteratively selects the most relevant dictionary elements to approximate the channel. The iterations continue until the residual error falls below a predefined threshold or the sparsity constraint is met. The implementation details of the OMP algorithm and the design of the training hybrid beamformer follow the principles outlined in [36], [37]. For brevity, these details are omitted here.